# SplitOut: Out-of-the-Box Training-Hijacking Detection in Split Learning via Outlier Detection

### v2

@@ 4.5 Computational Cost
As stated earlier, the core assumption behind SplitOut is that the clients can reasonably perform forward/backward passes over the entire model for a very small number of batches (e.g. 9 in case of F-/MNIST and 7 in CIFAR10/100). It has also been demonstrated as part of the earlier data reconstruction attacks [14, 36] that higher number of client-side layers makes client outputs harder to invert, giving rise to the privacy/cost trade-off inherent in many privacy-preserving systems. Thus, unlike earlier work on detecting training-hijacking attacks [13, 16] that relied on heuristical rather than computational assumptions as outlined in Section 2.4, our work extends this privacy/cost trade-off by instead giving up on task-related assumptions. This leads in turn to a simpler-to-use detection method that can be applied out-of-the-box without any hyperparameter tuning.
@@ 5 Related Work
Split learning has been shown to be vulnerable to various kinds of attacks beyond the training-hijacking literature we have covered. Regarding data privacy, split learning is vulnerable to model inversion attacks [14, 17, 41] aiming to reconstruct private client data, and label leakage attacks [14, 34, 33, 25], where the attacker (server) attempts to infer the private labels from the gradients sent by the client. Furthermore, confidentiality of the client models can be violated through model stealing attacks [14, 17], and the integrity of the system can be damaged through backdoor attacks [3].
Existence of such attacks has motivated the search for defense mechanisms. Differential privacy [12] has been applied to make private data reconstruction [18, 59] and label leakage [63] harder. Further work has considered integrating an additional optimization step to minimize the information the intermediate outputs leak [54], utilizing the permutation-equivariance of certain deep neural network components by shuffling the inputs without a cost to the model performance [61, 64], and using homomorphic encryption to compute part of the system on encrypted data to obtain formal guarantees on information leakage [26, 40].
@@ 6 Conclusion and Future Work
We demonstrated that SplitOut, an out-of-the-box outlier detection approach using the local outlier factor algorithm, can detect whether a split learning server is launching a training-hijacking attack or not with almost-perfect accuracy. Compared to earlier such detection work, we make stronger assumptions on the clients’ compute capabilities, but require fewer hyperparameters to be tuned, leading to a simpler to method that can be implemented readily using popular frameworks such as scikit-learn [39] on top of an existing split learning system. Due to the proactive design of our solution (performing outlier detection against honestly trained gradients), it can potentially detect some not-yet-known attacks.
Yet, a main limitation of SplitOut is that the space of training-hijacking attacks so far has been limited to feature-space alignment attacks. While such attacks have demonstrated strong performance for data reconstruction and implementing backdoors, future attacks employing a different training-hijacking approach could potentially behave in a different way than the observations we made in Section 4.4 and make our approach less effective. We leave this as future work.
Appendix A Comprehensive Results For Attack Detection
In this section, we provide our ablation study results. Clients collect training data for LOF by training the entire model over one epoch, with some portion of their data. During this training process, the impact of using their data at different fractions on attack detection was examined. SplitOut’s performance against FSHA [36], backdoor [65] and SplitSpy [16] is shown respectively in Tables I(a), II(a) and III(a), when trained with different shares of client training data (1%, 10%, 25%, 50%, 75%, 100%) and different decision-window sizes (1, 10).
TABLE I: Detecting FSHA using SplitOut [36]. True and false positive rates for our method averaged over 100 runs. LOF data rate is the percentage of client data used in simulating the entire model to obtain training data for the LOF algorithm. Window size is the number of most recent single-gradient decisions to perform a majority vote over to reach the final decision. The $t$ field denotes the average point of detection (as the share of total number of batches).
@@ Appendix A Comprehensive Results For Attack Detection
Dataset normal-$\rightarrow$ MNIST F-MNIST CIFAR10 CIFAR100
LOF Data Rate (%) normal-$\downarrow$ TPR FPR $t$ TPR FPR $t$ TPR FPR $t$ TPR FPR $t$
1 1.0 0.03 0.0013 1.0 0.02 0.0023 1.0 0.58 0.0043 1.0 0.70 0.0590
10 1.0 0.03 0.0013 1.0 0.01 0.0023 1.0 0.07 0.0076 1.0 0.23 0.0111
25 1.0 0.03 0.0013 1.0 0.01 0.0023 1.0 0.01 0.0089 1.0 0.10 0.0129
50 1.0 0.03 0.0013 1.0 0.01 0.0023 1.0 0.01 0.0097 1.0 0.02 0.0195
75 1.0 0.03 0.0013 1.0 0.01 0.0024 1.0 0.01 0.0108 1.0 0.03 0.0265
100 1.0 0.03 0.0013 1.0 0.00 0.0024 1.0 0.01 0.0117 1.0 0.02 0.0354
@@
Dataset normal-$\rightarrow$ MNIST F-MNIST CIFAR10 CIFAR100
LOF Data Rate (%) normal-$\downarrow$ TPR FPR $t$ TPR FPR $t$ TPR FPR $t$ TPR FPR $t$
1 1.0 0.0 0.0107 1.0 0.0 0.0109 1.0 0.02 0.0132 1.0 0.02 0.0142
10 1.0 0.0 0.0107 1.0 0.0 0.0109 1.0 0.00 0.0145 1.0 0.00 0.0219
25 1.0 0.0 0.0107 1.0 0.0 0.0109 1.0 0.00 0.0156 1.0 0.00 0.0289
50 1.0 0.0 0.0107 1.0 0.0 0.0109 1.0 0.00 0.0170 1.0 0.00 0.0357
75 1.0 0.0 0.0107 1.0 0.0 0.0109 0.99 0.00 0.0170 1.0 0.00 0.0434
100 1.0 0.0 0.0107 1.0 0.0 0.0109 0.99 0.00 0.0183 1.0 0.00 0.0557
TABLE II: Detecting the Backdoor Attack using SplitOut [65]. True and false positive rates for our method averaged over 100 runs. LOF data rate is the percentage of client data used in simulating the entire model to obtain training data for the LOF algorithm. Window size is the number of most recent single-gradient decisions to perform a majority vote over to reach the final decision. The $t$ field denotes the average point of detection (as the share of total number of batches).
@@
Dataset normal-$\rightarrow$ MNIST F-MNIST CIFAR10 CIFAR100
LOF Data Rate (%) normal-$\downarrow$ TPR FPR $t$ TPR FPR $t$ TPR FPR $t$ TPR FPR $t$
1 1.0 0.03 0.0011 1.0 0.02 0.0011 1.0 0.58 0.0026 1.0 0.70 0.0028
10 1.0 0.03 0.0011 1.0 0.01 0.0011 1.0 0.07 0.0036 1.0 0.23 0.0039
25 1.0 0.03 0.0011 1.0 0.01 0.0011 1.0 0.01 0.0038 1.0 0.10 0.0041
50 1.0 0.03 0.0011 1.0 0.01 0.0011 1.0 0.01 0.0039 1.0 0.02 0.0044
75 1.0 0.03 0.0011 1.0 0.01 0.0011 1.0 0.01 0.0039 1.0 0.03 0.0045
100 1.0 0.03 0.0011 1.0 0.00 0.0011 1.0 0.01 0.0040 1.0 0.02 0.0046
@@
Dataset normal-$\rightarrow$ MNIST F-MNIST CIFAR10 CIFAR100
LOF Data Rate (%) normal-$\downarrow$ TPR FPR $t$ TPR FPR $t$ TPR FPR $t$ TPR FPR $t$
1 1.0 0.0 0.0106 1.0 0.0 0.0107 1.0 0.02 0.0128 1.0 0.02 0.0128
10 1.0 0.0 0.0106 1.0 0.0 0.0107 1.0 0.00 0.0128 1.0 0.00 0.0128
25 1.0 0.0 0.0106 1.0 0.0 0.0107 1.0 0.00 0.0128 1.0 0.00 0.0128
50 1.0 0.0 0.0106 1.0 0.0 0.0107 1.0 0.00 0.0128 1.0 0.00 0.0128
75 1.0 0.0 0.0107 1.0 0.0 0.0107 1.0 0.00 0.0128 1.0 0.00 0.0128
100 1.0 0.0 0.0107 1.0 0.0 0.0107 1.0 0.00 0.0128 1.0 0.00 0.0128
TABLE III: Detecting SplitSpy using SplitOut [16]. True and false positive rates for our method averaged over 100 runs. LOF data rate is the percentage of client data used in simulating the entire model to obtain training data for the LOF algorithm. Window size is the number of most recent single-gradient decisions to perform a majority vote over to reach the final decision. The $t$ field denotes the average point of detection (as the share of total number of batches).
@@
Dataset normal-$\rightarrow$ MNIST F-MNIST CIFAR10 CIFAR100
LOF Data Rate (%) normal-$\downarrow$ TPR FPR $t$ TPR FPR $t$ TPR FPR $t$ TPR FPR $t$
1 1.0 0.03 0.001 1.0 0.02 0.0011 1.0 0.58 0.0052 1.0 0.70 0.0056
10 1.0 0.03 0.001 1.0 0.01 0.0011 1.0 0.07 0.0082 1.0 0.23 0.0101
25 1.0 0.03 0.001 1.0 0.01 0.0011 1.0 0.01 0.0100 1.0 0.10 0.1230
50 1.0 0.03 0.001 1.0 0.01 0.0011 1.0 0.01 0.0106 1.0 0.02 0.0171
75 1.0 0.03 0.001 1.0 0.01 0.0012 1.0 0.01 0.0102 1.0 0.03 0.0208
100 1.0 0.03 0.001 1.0 0.00 0.0012 1.0 0.01 0.0114 1.0 0.02 0.0227
@@
Dataset normal-$\rightarrow$ MNIST F-MNIST CIFAR10 CIFAR100
LOF Data Rate (%) normal-$\downarrow$ TPR FPR $t$ TPR FPR $t$ TPR FPR $t$ TPR FPR $t$
1 1.0 0.0 0.0106 1.0 0.0 0.0107 1.00 0.02 0.0137 1.0 0.02 0.0173
10 1.0 0.0 0.0106 1.0 0.0 0.0107 1.00 0.00 0.0100 1.0 0.00 0.0199
25 1.0 0.0 0.0106 1.0 0.0 0.0107 0.99 0.00 0.0158 1.0 0.00 0.0250
50 1.0 0.0 0.0106 1.0 0.0 0.0107 0.99 0.00 0.0164 1.0 0.00 0.0347
75 1.0 0.0 0.0106 1.0 0.0 0.0107 1.00 0.00 0.0169 1.0 0.00 0.0423
100 1.0 0.0 0.0106 1.0 0.0 0.0107 0.99 0.00 0.0174 1.0 0.00 0.0456

### v3

@@ 4.5 Computational Cost
As stated earlier, the core assumption behind SplitOut is that the clients can reasonably perform forward/backward passes over the entire model for a very small number of batches (e.g. 9 in case of F-/MNIST and 7 in CIFAR10/100). It has also been demonstrated as part of the earlier data reconstruction attacks [erdogan2021unsplit, pasquini_unleashing_2021] that higher number of client-side layers makes client outputs harder to invert, giving rise to the privacy/cost trade-off inherent in many privacy-preserving systems. Thus, unlike earlier work on detecting training-hijacking attacks [splitguard, fuFocusingPinocchioNose2023] that relied on heuristical rather than computational assumptions as outlined in Section 2.4, our work extends this privacy/cost trade-off by instead giving up on task-related assumptions. This leads in turn to a simpler-to-use detection method that can be applied out-of-the-box without any hyperparameter tuning.
@@ 5 Related Work
Split learning has been shown to be vulnerable to various kinds of attacks beyond the training-hijacking literature we have covered. Regarding data privacy, split learning is vulnerable to model inversion attacks [erdogan2021unsplit, gao2023pcat, qiuEXACTExtensiveAttack2023, zhu2023passive] aiming to reconstruct private client data, and label leakage attacks [erdogan2021unsplit, liu2022clustering, li2021label, kariyappa2023], where the attacker (server) attempts to infer the private labels from the gradients sent by the client. Furthermore, confidentiality of the client models can be violated through model stealing attacks [erdogan2021unsplit, gao2023pcat], and the integrity of the system can be damaged through backdoor attacks [bai2023villain].
Existence of such attacks has motivated the search for defense mechanisms. Differential privacy [dwork2014algorithmic] has been applied to make private data reconstruction [gawron2022feature, wu2023split] and label leakage [yang2022differentially] harder. Further work has considered integrating an additional optimization step to minimize the information the intermediate outputs leak [vepakomma_nopeek_2020], utilizing the permutation-equivariance of certain deep neural network components by shuffling the inputs without a cost to the model performance [xuShuffledTransformerPrivacyPreserving2023], and using homomorphic encryption to compute part of the system on encrypted data to obtain formal guarantees on information leakage [khan2023love].
@@ 6 Conclusion and Future Work
We demonstrated that SplitOut, an out-of-the-box outlier detection approach using the local outlier factor algorithm, can detect whether a split learning server is launching a training-hijacking attack or not with almost-perfect accuracy. Compared to earlier such detection work, we make stronger assumptions on the clients’ compute capabilities, but require fewer hyperparameters to be tuned, leading to a simpler to method that can be implemented readily using popular frameworks such as scikit-learn [scikit-learn] on top of an existing split learning system. Due to the proactive design of our solution (performing outlier detection against honestly trained gradients), it can potentially detect some not-yet-known attacks.
Limitations. A main limitation of SplitOut is that the space of training-hijacking attacks so far has been limited to feature-space alignment attacks. While such attacks have demonstrated strong performance for data reconstruction and implementing backdoors, future attacks employing a different training-hijacking approach could potentially behave in a different way and make our approach less effective. Another limitation of SplitOut is its applicability when the attacks start at a later epoch after some amount of honest training. While all the attacks we have considered start with the beginning of training and their generalizability against client models of various convergence levels is not trivial, it is nevertheless reasonable that honest gradients from a later epoch are classified as outliers by LOF trained with honest gradients from the first epoch. Since the previous detection methods [splitguard, fuFocusingPinocchioNose2023] also have this limitation, we leave this as future work to potentially improve the field.
Acknowledgements
We acknowledge the Scientific and Technological Research Council of Turkey (TÜBİTAK) project 119E088.
[…]
Appendix 0.A Attack Detection with One-class SVM
Table 1: Detecting FSHA using One-class SVM [pasquini_unleashing_2021]. True and false positive rates for our method averaged over 100 runs. One-class SVM data rate is the percentage of client data used in simulating the entire model to obtain training data for the one-class SVM algorithm. Window size is the number of most recent single-gradient decisions to perform a majority vote over to reach the final decision. The $t$ field denotes the average point of detection (as the share of the total number of batches), and the standard error of the mean is shown. Parameter nu was chosen as 0.0001, kernel was chosen as ‘rbf’, and gamma was chosen as auto’ in one-class SVM model training.
@@ Appendix 0.A Attack Detection with One-class SVM
Data Rate (%) TPR t FPR 1 1.0 $0.0013\pm 0.0$ 1.00 10 1.0 $0.0013\pm 0.0$ 1.00 25 1.0 $0.0013\pm 0.0$ 1.00 50 1.0 $0.0013\pm 0.0$ 1.00 75 1.0 $0.0018\pm 0.0002$ 1.00 100 1.0 $0.0023\pm 0.0002$ 0.98
@@
Data Rate (%) TPR t FPR 1 1.0 $0.0128\pm 0.0$ 1.00 10 1.0 $0.0128\pm 0.0$ 1.00 25 1.0 $0.0128\pm 0.0$ 1.00 50 1.0 $0.0128\pm 0.0$ 1.00 75 1.0 $0.0128\pm 0.0$ 0.02 100 1.0 $0.0129\pm 0.0$ 0.00
We experimented with one-class SVM [scholkopf1999support] as another outlier detection model. However, one-class SVM, which is based on applying a prediction method to new data with a decision boundary, is sensitive to outliers and it can be shown as one of its few major disadvantages compared to LOF. Additionally, the number of parameters that need to be determined to train a successful one-class SVM model is considerably more compared to LOF, while LOF only needs the number of neighbors for training.
In our experiments, we performed a hyperparameter search for kernel, gamma and nu values to train the one-class SVM model, which we expected to keep the FPR value low while keeping the TPR value high. Since one-class SVM tries to draw boundaries for the existing data, and the kernel parameter determines the kernel type to be used in the algorithm. Another parameter, gamma coefficient should be selected for ’rbf’, ’poly’ and ’sigmoid’ kernels. Parameter nu defines the upper bound of the ratio of outlier or training errors and the lower bound of the ratio of support vectors. Valid range of nu for one-class SVM training should be (0,1].
We measured the performance of one-class SVM for FSHA [pasquini_unleashing_2021] on CIFAR10 dataset with values for the kernel parameter (linear, poly, rbf, sigmoid, precomputed), for the gamma parameter (scale, auto), and for the nu parameter (0.0001, 0.001, 0.01, 0.1, 0.5).
As a result of our experiments, we observed that the most successful results were obtained when nu was 0.0001, kernel was ‘rbf’ and gamma parameter was ’auto’. As can be seen in Table 1, we observed that one-class SVM can approach the results of LOF only when the data rate is high (75%, 100%), while LOF can give the same results with low data rates such as 1% and 10%.
Table 2: Detecting FSHA [pasquini_unleashing_2021] using SplitOut for EMNIST-digits dataset. True and false positive rates for our method averaged over 50 runs. The $t$ field denotes the average point of detection (as the share of the total number of batches), and the standard error of the mean is shown.
@@
Data Rate (%) TPR $t$ FPR 0.1 1 $4.1x10^{-6}\pm 0.0$ 1 0.25 1 $4.1x10^{-6}\pm 0.0$98 1 1 $4.1x10^{-6}\pm 0.0$98 10 1 $4.1x10^{-6}\pm 0.0$94 25 1 $5.0x10^{-6}\pm 5.4x10^{-7}$40 50 1 $9.9x10^{-6}\pm 1.3x10^{-6}$04 75 1 $1.4x10^{-5}\pm 1.6x10^{-6}$ 0 100 1 $1.6x10^{-5}\pm 1.6x10^{-6}$ 0
@@
Data Rate (%) TPR $t$ FPR 0.1 1 $4.1x10^{-5}\pm 0.0$ 0 0.25 1 $4.1x10^{-5}\pm 0.0$ 0 1 1 $4.1x10^{-5}\pm 0.0$ 0 10 1 $4.1x10^{-5}\pm 0.0$ 0 25 1 $4.2x10^{-5}\pm 1.2x10^{-7}$ 0 50 1 $4.3x10^{-5}\pm 4.2x10^{-7}$ 0 75 1 $4.3x10^{-5}\pm 7.4x10^{-7}$ 0 100 1 $4.4x10^{-5}\pm 7.4x10^{-7}$ 0
Appendix 0.B Comprehensive Results For Attack Detection
In this section, we provide our ablation study results. Clients collect training data for LOF by training the entire model over one epoch, with some portion of their data. During this training process, the impact of using their data at different fractions on attack detection was examined. SplitOut’s performance against FSHA [pasquini_unleashing_2021], backdoor [yuHowBackdoorSplit2023] and SplitSpy [fuFocusingPinocchioNose2023] is shown respectively in Tables 5, 6, and 4 when trained with different shares of client training data (1%, 10%, 25%, 50%, 75%, 100%) and different decision-window sizes (1, 10). The LOF data rate is the percentage of client data used in simulating the entire model to obtain training data for the LOF algorithm. Window size is the number of most recent single-gradient decisions to perform a majority vote over to reach the final decision.
The performance of Split Out against FSHA was also experimented for EMNIST-digits dataset. The results for the EMNIST-digits dataset containing 3750 batches with batch size 64 as used in all setups are shown in Table 2.
Appendix 0.C Comprehensive Results For Attack Detection in Later Epochs
To evaluate SplitOut’s performance in later epochs, Table 3 includes detection statistics for when the attack starts in the second epoch. In this case, it is possible to completely detect the attack (with 0 FPR) with the SplitOut where the LOF model is trained with the first epoch gradients, although more data is needed than had the attack started in the first epoch, as can be seen in Table 3. Besides, in Figure 1, we added t-SNE plots to show the randomness and similarity properties of the honest and malicious gradients from the 2nd epoch and the 1st epoch gradients.
Although previously published defense methods [splitguard, fuFocusingPinocchioNose2023] against training-hi-jacking have been created by considering the split learning protocol between the client and the honest server in cases where the attack is initiated simultaneously, the attacks do not strictly require starting against randomly-initialized clients. It is possible that an adaptive attack method can start in later epochs, but there is a trade-off for the attacker who trades attack time for more difficult detectability. Nevertheless, the attacks presented so far have started from the beginning of training, and their generalizability against client models of various convergence levels on different tasks is not trivial.
Figure 1: T-SNE [van2008visualizing] dimension reduction comparing honest and malicious (FSHA) gradients obtained from a randomly chosen run of the first and second epoch of training on CIFAR10.
Table 3: Detecting adaptive FSHA (in later epochs) using SplitOut [pasquini_unleashing_2021]. True and false positive rates for our method averaged over 50 runs. The $t$ field denotes the average point of detection (as the share of the total number of batches), and the standard error of the mean is shown.
@@ Appendix 0.C Comprehensive Results For Attack Detection in Later Epochs
Method $\rightarrow$ 1st Epoch Training 2nd Epoch Training 1st+2nd Epoch Training Data % $\downarrow$ TPR FPR t TPR FPR t TPR FPR t 1 1 1 .$0013\pm 0$ 1 .64 .$0076\pm.0014$ 1 1 .$0013\pm.0000$ 10 1 1 .$0013\pm 0$94 .06 .$0231\pm.0045$ 1 1 .$0013\pm.0001$ 25 1 1 .$0013\pm.0001$92 .02 .$0309\pm.0056$ 1 .92 .$0035\pm.0009$ 50 1 .92 .$0035\pm.0009$88 .02 .$0346\pm.0063$96 .10 .$0212\pm.0047$ 75 .98 .36 .$0130\pm.0028$82 .00 .$0886\pm.0278$92 .02 .$0356\pm.0068$ 100 .96 .10 .$0212\pm.0047$78 .00 .$0970\pm.0291$80 .00 .$0932\pm.0285$
@@
Method $\rightarrow$ 1st Epoch Training 2nd Epoch Training 1st+2nd Epoch Training Data % $\downarrow$ TPR FPR t TPR FPR t TPR FPR t 1 1 1 .$0013\pm 0$96 0 .$0452\pm.0093$ 1 1 .$00128\pm 0$ 10 1 1 .$0013\pm 0$68 0 .$0915\pm.0182$ 1 .84 .$00128\pm 0$ 25 1 .68 .$0129\pm.0002$44 0 .$1179\pm.0257$ 1 .00 .$0025\pm.0054$ 50 1 .00 .$0250\pm.0054$38 0 .$1623\pm.0445$78 .00 .$1042\pm.0178$ 75 .88 .00 .$0534\pm.0104$28 0 .$1568\pm.0524$40 .00 .$1153\pm.0260$ 100 .78 .00 .$1042\pm.0178$24 0 .$1163\pm.0233$28 .00 .$1595\pm.0520$
Table 4: Detecting FSHA using SplitOut [pasquini_unleashing_2021]. True and false positive rates for our method averaged over 100 runs. The $t$ field denotes the average point of detection (as the share of the total number of batches), and the standard error of the mean is shown.
(a) Window size: 1
Dataset $\rightarrow$ MNIST F-MNIST CIFAR10 CIFAR100 Data % $\downarrow$   TPR FPR $t$   TPR FPR $t$   TPR FPR $t$   TPR FPR $t$ 1 1 0 .$0013\pm 0$ 1 0 .$0023\pm.0002$ 1 .98 .$0019\pm.0002$ 1 .98 .$0022\pm.0002$ 10 1 0 .$0013\pm 0$ 1 0 .$0025\pm.0002$ 1 .96 .$0019\pm.0002$ 1 .98 .$0019\pm.0002$ 25 1 0 .$0013\pm 0$ 1 0 .$0025\pm.0002$ 1 .52 .$0044\pm.0003$ 1 .97 .$0027\pm.0002$ 50 1 0 .$0013\pm 0$ 1 0 .$0025\pm.0002$ 1 .03 .$0078\pm.0002$ 1 .43 .$0072\pm.0003$ 75 1 0 .$0013\pm 0$ 1 0 .$0025\pm.0002$ 1 .00 .$0093\pm.0003$ 1 .04 .$0107\pm.0005$ 100 1 0 .$0013\pm 0$ 1 0 .$0025\pm.0002$ 1 .00 .$0186\pm.0011$ 1 .00 .$0283\pm.0061$
(b) Window size: 10
@@
Dataset $\rightarrow$ MNIST F-MNIST CIFAR10 CIFAR100 Data % $\downarrow$   TPR FPR $t$   TPR FPR $t$   TPR FPR $t$   TPR FPR $t$ 1 1 0 .$0107\pm 0$ 1 0 .$0109\pm.0001$ 1 .02 .$0128\pm 0$ 1 .14 .$0129\pm 0$ 10 1 0 .$0107\pm 0$ 1 0 .$0109\pm.0001$ 1 .03 .$0128\pm 0$ 1 .41 .$0129\pm 0$ 25 1 0 .$0107\pm 0$ 1 0 .$0109\pm.0001$ 1 .00 .$0132\pm.0001$ 1 .14 .$0129\pm 0$ 50 1 0 .$0107\pm 0$ 1 0 .$0109\pm.0001$ 1 .00 .$0145\pm.0002$ 1 .00 .$0145\pm.0002$ 75 1 0 .$0107\pm 0$ 1 0 .$0109\pm.0001$ 1 .00 .$0161\pm.0003$ 1 .00 .$0190\pm.0011$ 100 1 0 .$0107\pm 0$ 1 0 .$0109\pm.0001$99 .00 .$0186\pm.0011$ 1 .00 .0808±.0081plus-or-minus.0808.0081.$08$0081
Table 5: Detecting the Backdoor Attack using SplitOut [yuHowBackdoorSplit2023]. True and false positive rates for our method averaged over 100 runs. The $t$ field denotes the average point of detection (as the share of the total number of batches), and the standard error of the mean is shown.
(a) Window size: 1
Dataset $\rightarrow$ MNIST F-MNIST CIFAR10 CIFAR100 Data % $\downarrow$   TPR FPR $t$   TPR FPR $t$   TPR FPR $t$   TPR FPR $t$ 1 1 0 .$0011\pm 0$ 1 0 .$0011\pm 0$ 1 .98 .$0018\pm.0001$ 1 .98 .$0014\pm.0001$ 10 1 0 .$0011\pm 0$ 1 0 .$0011\pm 0$ 1 .96 .$0018\pm.0001$ 1 .98 .$0014\pm.0001$ 25 1 0 .$0011\pm 0$ 1 0 .$0011\pm 0$ 1 .52 .$0026\pm.0001$ 1 .97 .$0015\pm.0001$ 50 1 0 .$0011\pm 0$ 1 0 .$0011\pm 0$ 1 .03 .$0034\pm.0001$ 1 .43 .$0051\pm.0003$ 75 1 0 .$0011\pm 0$ 1 0 .$0011\pm 0$ 1 .00 .$0038\pm.0000$ 1 .04 .$0083\pm.0004$ 100 1 0 .$0011\pm 0$ 1 0 .$0011\pm 0$ 1 .00 .$0040\pm.0001$ 1 .00 .$0176\pm.0043$
(b) Window size: 10
Dataset $\rightarrow$ MNIST F-MNIST CIFAR10 CIFAR100 Data % $\downarrow$   TPR FPR $t$   TPR FPR $t$   TPR FPR $t$   TPR FPR $t$ 1 1 0 .$0107\pm 0$ 1 0 .$0107\pm 0$ 1 .02 .$0128\pm 0$ 1 .14 .$0128\pm 0$ 10 1 0 .$0107\pm 0$ 1 0 .$0107\pm 0$ 1 .03 .$0128\pm 0$ 1 .41 .$0128\pm 0$ 25 1 0 .$0107\pm 0$ 1 0 .$0107\pm 0$ 1 .00 .$0128\pm 0$ 1 .14 .$0128\pm 0$ 50 1 0 .$0107\pm 0$ 1 0 .$0107\pm 0$ 1 .00 .$0128\pm 0$ 1 .00 .$0134\pm.0001$ 75 1 0 .$0107\pm 0$ 1 0 .$0107\pm 0$ 1 .00 .$0128\pm 0$ 1 .00 .$0173\pm.0019$ 100 1 0 .$0107\pm 0$ 1 0 .$0107\pm 0$ 1 .00 .$0128\pm 0$ 1 .00 .$0308\pm.0043$
Table 6: Detecting SplitSpy using SplitOut [fuFocusingPinocchioNose2023]. True and false positive rates for our method averaged over 100 runs. The $t$ field denotes the average point of detection (as the share of the total number of batches), and the standard error of the mean is shown.
(a) Window size: 1
Dataset $\rightarrow$ MNIST F-MNIST CIFAR10 CIFAR100 Data % $\downarrow$   TPR FPR $t$   TPR FPR $t$   TPR FPR $t$   TPR FPR $t$ 1 1 0 .0011$\pm$0 1 0 .0012$\pm$.0001 1 .98 .0021$\pm$.0002 1 .98 .0023$\pm$.0002 10 1 0 .0011$\pm$0 1 0 .0012$\pm$.0001 1 .96 .0020$\pm$.0002 1 .98 .0019$\pm$.0002 25 1 0 .0011$\pm$0 1 0 .0012$\pm$.0001 1 .52 .0049$\pm$.0003 1 .97 .0026$\pm$.0002 50 1 0 .0011$\pm$0 1 0 .0012$\pm$.0001 1 .03 .0079$\pm$.0003 1 .43 .0068$\pm$.0003 75 1 0 .0011$\pm$0 1 0 .0012$\pm$.0001 1 .00 .0096$\pm$.0003 1 .04 .0108$\pm$.0004 100 1 0 .0011$\pm$0 1 0 .0012$\pm$.0001 1 .00 .0115$\pm$.0008 1 .00 .0195$\pm$.0033
(b) Window size: 10
Dataset $\rightarrow$ MNIST F-MNIST CIFAR10 CIFAR100 Data % $\downarrow$   TPR FPR $t$   TPR FPR $t$   TPR FPR $t$   TPR FPR $t$ 1 1 0 .0107$\pm$0 1 0 .0107$\pm$0 1 .02 .0129$\pm$0 1 .14 .0129$\pm$0 10 1 0 .0107$\pm$0 1 0 .0107$\pm$0 1 .03 .0129$\pm$0 1 .41 .0129$\pm$0 25 1 0 .0107$\pm$0 1 0 .0107$\pm$0 1 .00 .0133$\pm$.0001 1 .14 .0129$\pm$0 50 1 0 .0107$\pm$0 1 0 .0107$\pm$0 1 .00 .0148$\pm$.0002 1 .00 .0144$\pm$.0002 75 1 0 .0107$\pm$0 1 0 .0107$\pm$0 1 .00 .0163$\pm$.0003 1 .00 .0206$\pm$.0029 100 1 0 .0107$\pm$0 1 0 .0107$\pm$0 .98 .00 .0176$\pm$.0004 1 .00 .0422$\pm$.0072
Table 7: Detecting FSHA using SplitOut [pasquini_unleashing_2021] where client and server have different models. True and false positive rates for our method averaged over 50 runs. The $t$ field denotes the average point of detection (as the share of the total number of batches), and the standard error of the mean is shown.
(a) Window size: 1
Dataset $\rightarrow$ MNIST F-MNIST CIFAR10 CIFAR100 Data % $\downarrow$   TPR FPR $t$   TPR FPR $t$   TPR FPR $t$   TPR FPR $t$ 1 1 .84 .0011$\pm$0 1 .64 .0011$\pm$0 1 1 .0015$\pm$.0002 1 1 .0014$\pm$.0001 10 1 .84 .0011$\pm$0 1 .80 .0011$\pm$0 1 1 .0015$\pm$.0002 1 1 .0014$\pm$.0001 25 1 .84 .0011$\pm$0 1 .80 .0011$\pm$0 1 1 .0029$\pm$.0004 1 1 .0016$\pm$.0002 50 1 .84 .0011$\pm$0 1 .82 .0011$\pm$0 1 .98 .0069$\pm$.0003 1 1 .0050$\pm$.0004 75 1 .84 .0011$\pm$0 1 .82 .0011$\pm$0 1 .88 .0083$\pm$.0002 1 .72 .0088$\pm$.0007 100 1 .84 .0011$\pm$0 1 .82 .0011$\pm$0 1 .76 .0092$\pm$.0004 1 .38 .0121$\pm$.0011
(b) Window size: 10
Dataset $\rightarrow$ MNIST F-MNIST CIFAR10 CIFAR100 Data % $\downarrow$   TPR FPR $t$   TPR FPR $t$   TPR FPR $t$   TPR FPR $t$ 1 1 0 .0107$\pm$0 1 0 .0107$\pm$0 1 1 .0128$\pm$0 1 1 .0128$\pm$0 10 1 0 .0107$\pm$0 1 0 .0107$\pm$0 1 .98 .0128$\pm$0 1 1 .0128$\pm$0 25 1 0 .0107$\pm$0 1 0 .0107$\pm$0 1 .76 .0130$\pm$.0001 1 1 .0128$\pm$.0000 50 1 0 .0107$\pm$0 1 0 .0107$\pm$0 1 .16 .0139$\pm$.0002 1 .50 .0134$\pm$.0001 75 1 0 .0107$\pm$0 1 0 .0107$\pm$0 1 .04 .0148$\pm$.0003 1 .06 .0197$\pm$.0038 100 1 0 .0107$\pm$0 1 0 .0107$\pm$0 1 .00 .0159$\pm$.0005 1 .02 .0227$\pm$.0038